OpenStudy (ujjwal):

let f(x) be a quadratic polynomial. if, f(0),f(1),f(-1)\(\in\mathbb{z}\), then show that f(n)\(\in\mathbb{z}\) for all \(n\in\mathbb{z}\)

OpenStudy (anonymous):

By PMI ?

OpenStudy (ujjwal):

what's that?

OpenStudy (anonymous):

Principle of Mathametical Induction

OpenStudy (anonymous):

let f(x) = ax^2 + bx + c

OpenStudy (ujjwal):

Sadly, mathematical induction holds only for natural numbers, not all integers..

OpenStudy (amistre64):

let n1,n2,n3 be some integers f(x) = n1 + n2x + n3x(x+1) f(0) = n1 f(-1) = n1 - n2 f(1) = n1 + n2 + 2n3 satisfies the conditions

OpenStudy (amistre64):

i spose n was a bad choice for genericy ...

OpenStudy (amistre64):

show that for all z, z n2 and z(z+1)n3 are integers and closed

OpenStudy (amistre64):

or in other words: Z is closed under addition and multiplication therefore z n2 is an integer; and z(z+1) n3 is an integer

OpenStudy (amistre64):

let a,b,c,n in Z; and let f(x) = a + bx + cx(x+1) be a quadratic polynomial f(0) = a is an integer f(-1) = a - b is an integer since Z is closed under addition/subtraction f(1) = a + b + 2c is an integer since Z is closed under addition and multipication let x = n f(n) = a + bn + c n(n+1) is an integer since Z is closed .....

OpenStudy (ujjwal):

Does, this not only prove for integral quadratic polynomial? can't there be other cases too where a,b,c are not integers, yet f(0), f(1), f(-1) are integers?

OpenStudy (amistre64):

thats not what it is asking for .....

OpenStudy (amistre64):

let f(x) be a quadratic poly ... NOT, let f(x) be all quadratic polys

OpenStudy (ujjwal):

I understand your solution but here you are taking an integral quadratic question which is just a special case of quadratic equation.

OpenStudy (amistre64):

spose we introduce a real number, d such that f(x) = a + bx + cx(x+1) + dx(x+1)(x-1) then for f(-1,0,1) we have integer solutons; but we cannot guarentee integer solutions for any other "n" in Z

OpenStudy (amistre64):

i am taking a quadratic polynomial that has integer solutions for x= -1,0,1

OpenStudy (amistre64):

lol, i turned d into a cubic

OpenStudy (ujjwal):

this can still be proved by using equation ax^2 +bx +c can't it be?

OpenStudy (amistre64):

i see what your trying to say tho; you think that my f(x) is not general enough

OpenStudy (amistre64):

i do not believe that it can be proved in that manner ... at least nothing comes to mind c would have to be an integer and a and b would still have to be integers when x=1 and x=-1; unless you can find some non integer values that add up to an integer (some fractions perhaps, but not all fractions work)

OpenStudy (ujjwal):

ok, supposing a,b,c are integers and f(x)=ax^2+bx+c we have f(0)=c f(1)=a+b+c f(-1)=a-b+c which are all integers and again, n is an integer, so, f(n)=an^2+bn+c is an integer Is there any special reason why you choose that particular equation?

OpenStudy (amistre64):

if a and b are fractions; then they would have to satisfy some b = z-a relationship maybe p/q n^2 + s/t n n (p/q n + s/t) p/q n + s/t = z p/q = (z - s/t)/n

OpenStudy (amistre64):

i simply built it so that each term is zeroed out is all

OpenStudy (ujjwal):

So, it holds true for any arbitrary quadratic equation as long as all co-efficients are integers?

OpenStudy (amistre64):

even in the case where you could use rational numbers for a and b there is still a limited set of fraction that would be available to use and therefore not be "generic" enough to prove ALL poly quads that have f(-1,0,1) in Z

OpenStudy (amistre64):

yes, f(n) = an^2 + bn + c; would by default have all integer solutions if abc in Z

OpenStudy (amistre64):

\[[f(n)=\frac pqn^2+\frac stn+c]\in Z,~if~(\frac pq=\frac{k-(s/t)}{n})~for~k\in Z, n\ne0\]

OpenStudy (ujjwal):

I guess, i will go with the above solution! :D and stop thinking about fractions..

OpenStudy (amistre64):

to find all possible cases: let a,b,c in Z such that 0 0 1 a 1 1 1 b 1 -1 1 c row reduces to give us something like:\[f(x)=\frac{b+c-2a}{2}x^2+\frac{b-c}{2}x+a\]

OpenStudy (ujjwal):

This one seems more complicated..

OpenStudy (amistre64):

it is, since we are considering all cases of f(x) that would give us integers for f(-1,0,1); if for all bc in Z this pans out then we have a proof

OpenStudy (amistre64):

if: \[\frac{b+c-2a}{2}n+\frac{b-c}{2}=k~:~k\in Z\] we will have our proof

OpenStudy (ujjwal):

why do we need to find such cases? It is clearly mentioned in question that f(0), f(1), f(-1) are integers..

OpenStudy (ujjwal):

and when they are integers, we have to prove f(x) is an integer too for all x belongs to R

OpenStudy (amistre64):

yes, it says that the results are integers; so letting abc be integer solutions we are able to form a matrix that we can row reduce to determine all suitable values for the coefficients that will produce the results

OpenStudy (ujjwal):

Z*

OpenStudy (amistre64):

since a quad poly is uniquely defined by 3 points, we are able to construct all possible f(-1,0,1) results in this manner

OpenStudy (ujjwal):

I don't think this question is that complicated.. May be the above method is sufficient! Thanks!

OpenStudy (amistre64):

\[\frac{b+c-2a}{2}n+\frac{b-c}{2}=k\] \[n(b+c-2a)+b-c=2k\] \[b(n+1)+c(n-1)-2n~a=2k\] since all we have remaining are integer products and sums, that should be proofed